# A Multi-Modality Evaluation of the Reality Gap in Autonomous Driving Systems

### v1

@@ Abstract
Simulation-based testing is a cornerstone of Autonomous Driving System (ADS) development, offering safe and scalable evaluation across diverse driving scenarios. However, discrepancies between simulated and real-world behavior—known as the reality gap—challenge the transferability of test results to deployed systems. In this paper, we present a comprehensive empirical study comparing four representative testing modalities: Software-in-the-Loop (SiL), Vehicle-in-the-Loop (ViL), Mixed-Reality (MR), and full real-world testing. Using a small-scale physical vehicle equipped with real sensors (camera and LiDAR), and its digital twin, we implement each setup and evaluate two ADS architectures (modular and end-to-end) across diverse indoor driving scenarios involving real obstacles, road topologies, and indoor environments. We systematically assess the impact of each testing modality along three dimensions of the reality gap: actuation, perception, and behavioral fidelity. Our results show that while SiL and ViL setups simplify critical aspects of real-world dynamics and sensing, MR testing improves perceptual realism without compromising safety or control. Importantly, we identify the conditions under which failures do not transfer across testing modalities and isolate the underlying dimensions of the gap responsible for these discrepancies. Our findings offer actionable insights into the respective strengths and limitations of each modality and outline a path toward more robust and transferable validation of autonomous driving systems.
@@ I Introduction
Various strategies have been proposed to mitigate the reality gap. Some aim to increase simulation fidelity through high-precision modeling (digital twins [9, 10, 11]), other address specific gap dimensions, such as the perception gap, by translating simulated sensor outputs into more realistic versions using generative models [12, 13, 14, 8, 15]. However, these methods are limited to model-level testing [16, 17] and do not capture the system-level interactions between perception, planning, and control modules that govern vehicle motion. As a result, they are susceptible to actuation gaps and often miss critical system-level failures [16, 14, 18]. Other strategies involve vehicle-in-the-loop (ViL) and mixed reality (MR) testing [19], by integrating physical components such as ECUs, small-scale robots, or full vehicles-into simulation loops. While ViL provides closed-loop evaluation, it remains partially virtualized and fails to capture real-world imperfections, such as sensor noise and lighting variability (perception gap) [20]. MR partially mitigates this by injecting virtual elements (e.g., obstacles) into real sensor data, enriching scenario realism. Prior system-level studies using small-scale robots investigating failure transferability [8, 21, 22, 23, 24, 25] have primarily documented the existence of the reality gap, without isolating their root causes or comparing how different test modalities influence gap reduction.
To this aim, in this paper, we conduct an empirical study of the reality gap in autonomous driving by comparing four representative testing modalities: SiL, ViL, MR, and full real-world execution (RW). Our goal is to characterize the dimensions of the reality gap-specifically along actuation, perception, and behavioral fidelity-and to assess the degree to which each testing setup retains ADS behavior relative to real-world ground truth behavior. While prior work has evaluated or mitigated specific aspects of the reality gap, a broader evaluation spanning multiple testing modalities and ADS architectures remains unaddressed.
@@ III-B1 Simulator
The Donkey Car framework provides a high-fidelity Unity3D [39] simulator that models the physical vehicle using the Nvidia PhysX engine [40]. We extended the simulator by developing a new Unity scene that procedurally generates track layouts using Catmull-Rom splines [41], based on real-world lane margin data from Room Nominal and Generalization. To replicate the physical environments, we applied high-resolution (48 MP) floor images as textures and configured the lane markings to match each sandbox. In addition, we modified the simulator to support a virtual depth sensor that mimics the behavior of the ToF system used in our real-world setup. The virtual sensor produces depth data at the same resolution and provides intrinsic matrix parameters consistent with those retrieved by ToF system.
@@ III-B2 Testing Framework
For the ViL and MR modalities, Figure 2 shows how the framework manages data flow between the simulator, the physical vehicle, and the ADS. Simulated data is first generated by the simulator. In the ViL setting, the full simulated sensor stream (e.g., camera or LiDAR) is used as input ❶. In the MR setting, only selected features or objects are rendered and extracted ❶(b), then merged with real sensor data from the physical vehicle ❷ to form mixed-reality inputs. The resulting sensor data-fully simulated in ViL or blended in MR-is processed by the ADS ❸, which outputs control commands for the physical vehicle ❹. The vehicle’s motion is tracked in real time ❺, and its pose is fed back to the simulator ❻ to maintain alignment between virtual and real environments. This forms a closed-loop system in which the ADS continuously perceives and acts on synchronized sensor data and vehicle dynamics.
@@ IV-D Experimental Setup Validation
For ViL and MR, the real vehicle’s motion must be continuously mapped into the simulation, requiring real-time execution to keep the simulated state and generated sensor data synchronized with the physical vehicle. Across all domains, the E2E or modular pipeline must execute fast enough to deliver control commands (simulated or real) without latency, as delays could distort the observed behavior. To verify this, we benchmarked the framework over 5000 samples, measuring execution time for every component. All modalities (SiL, ViL, MR, RW) ran in real time: the simulator achieved $\approx 80$ FPS ($\approx 12.5$ ms per frame) on our workstation (AMD Ryzen 5, 16 GB RAM, NVIDIA RTX 2060). The tracking system has been measured (at the simulator-level) to provide state updates at $\approx 99.93$ Hz ($\approx 10.01$ ms). Since the onboard camera and LiDAR sensors operate at 20 FPS (50 ms per frame), we configure the simulator to run its camera at 20 FPS and to avoid unnecessary system load, we cap the global simulation frame-rate at 60 FPS. The E2E ADS executed in $\approx 15$ ms on average, and the modular ADAS in $\approx 20$ ms. Inference begins as soon as a new sensor message arrives and can run in parallel if needed, though both pipelines consistently completed before the next frame. Control commands, measured at the vehicle-level, were issued every $\approx 50.5$ ms, remaining stable across all experiments.
@@ IV-E2 RQ2 (actuation gap)
To address RQ2, we quantify the actuation gap by comparing the physical response of the real vehicle to that of a simulated vehicle under identical control inputs. We limit the evaluation to SiL, ViL, and RW, omitting MR as this modality does not change actuation fidelity beyond what is already captured by ViL. Each experiment was repeated five times to capture variability, and data were logged at 100 Hz, yielding large frame-level sample sizes. Statistical analysis is applied only to per-frame metrics, where sample sizes support meaningful significance testing. We design five test scenarios spanning both low-level actuation and high-level control, described next. Forward Motion. Constant throttle is applied with steering fixed at zero and braking disabled. Throttle commands are taken directly from the RW-recorded ROS messages with matched timestamps. Trials run for three seconds using throttle values of 0.34, 0.365, and 0.39, which represent the lowest throttle at which the vehicle starts moving, to the highest throttle at which the vehicle can perform the experiment whilst remaining in the tracking area. We compute the total distance traveled, average speed, and trajectory, and compare both speed and trajectory distributions. This setup results in 15 executions (3 speed values × 5 repetitions).
@@ V-D1 Internal validity
We compared all ADS under identical parameter settings. One threat to internal validity concerns our custom implementation. However, this was unavoidable as no similar evaluation frameworks are available, to the best of our knowledge. Another threat may be due to our data collection phase and training of ADS, which may exhibit a large number of misbehaviors if trained inadequately or with poor quality data. We mitigated this threat by training and fine-tuning the best publicly available driving models, which performed consistently in nominal RW conditions. The RW-trained E2E model completed scenarios N1 and N2 in all five RW runs without failure (Table 1, Nominal–E2E–Real), indicating adequate training. The LiDAR-based modular pipeline failed in N2 due to late-appearing obstacles, as its perception relies on deterministic clustering rather than learned models; however, the ablation study shows the planning module operates correctly when decoupled from LiDAR inputs. Failures in scenario G stem from domain shift rather than poor training: the E2E ADS succeeds in SiL but fails in RW, which would not occur if the model were fundamentally incapable of solving the task. Overall, these results confirm that our comparisons are not confounded by under-trained or low-quality models.
@@ V-D2 External validity
We used a limited number of ADS architecture models in our evaluation, which we mitigated by covering representative ADS architectures. We considered only two physical tracks and a scaled-down platform, which may not capture all real-world physics (e.g., suspension dynamics or vehicle mass distribution). However, our goal is not to model absolute performance but to assess relative fidelity across modalities. Donkey Car was used as a proxy for full-size ADS also in previous studies [57, 21, 22, 23, 8, 14] and uniquely satisfies our requirements for studying transferability between simulated and real-world testing of ADS. Other platforms, such as DeepPiCar [58] and JetRacer [59], lack integrated simulators; Roboracer [60] offer only low-fidelity physical simulations that do not realistically capture real-world driving dynamics. AWS DeepRacer [61] is tightly integrated with AWS infrastructure and is primarily designed for reinforcement learning use cases, which are outside the scope of this work. In contrast, Donkey Car has been successfully adopted in numerous real-vehicle autonomous driving studies [57, 21, 22, 23, 8, 14, 36, 37, 38], making it a practical and cost-effective experimentation platform.
@@ VI-A Dissecting and Addressing the ADS Reality Gap
To mitigate the former in SiL, without relying on physical vehicles, recent neural rendering approaches have been proposed. Examples include generative AI for translating simulated into photorealistic images [12, 8, 14, 62, 13, 15] and diffusion models for generating realistic operational design domains [63, 64, 65]. Some are already integrated in simulators such as NVIDIA Omniverse [66]. While promising, these methods still suffer from correctness issues (e.g., artifacts, hallucinations), increase runtime cost, and leave the actuation gap and simulator unreliability [67, 68] unaddressed.
@@ VII-A Reality Gap Assessment Studies
The reality gap has been subjected to active research in many fields, including robotics, automotive, and artificial intelligence. For a comprehensive survey, we refer the reader to Hu et al. [78], while notable contributions are also provided in the software engineering community [1, 67, 8, 2, 3]. Concerning empirical studies on the reality gap, Stocco et al. [8] compare ADS lane-keeping models in simulated and physical environments, highlighting critical shortcomings that contribute to the gap. In this work, beyond assessing the gap between SiL and RW, we also evaluate mitigation strategies such as ViL and MR. Similarly, Gao et al.[79] propose MultiTest, a physically-aware object insertion framework for testing the robustness of fusion-based perception systems, while Gao et al. [80] outline key challenges in benchmarking AI-enabled multi-sensor fusion across diverse conditions. However, these studies focus primarily on perception robustness within simulated environments and do not extend to full-system evaluation. In contrast, our work provides, to the best of our knowledge, the first unified, system-level assessment of ViL and MR as reality gap mitigation strategies, explicitly analyzing their impact across perception, actuation, and behavior during live execution in SiL, MR, and RW settings under consistent and controlled conditions. However, the transferability to real-world vehicles is not assessed.
@@ VII-B Reality Gap Mitigation Studies
To address the perception gap, researchers have explored the use of generative AI techniques to translate simulated data into more realistic representations. These efforts include the generation of photorealistic images[12, 8, 14, 13, 15], realistic LiDAR point clouds [82], and methods for synthesizing operational design domains with high visual fidelity [63, 64, 83, 84]. While existing methods enhance perceptual realism, they typically run offline at the model level or in SiL, targeting single sensor modalities. In contrast, our work evaluates perception fidelity during live, system-level execution across SiL, MR, and RW, directly comparing how perceptual gaps affect the autonomous system system behavior.

### v2

@@ Abstract
Simulation-based testing is a cornerstone of Autonomous Driving System (ADS) development, offering safe and scalable evaluation across diverse driving scenarios. However, discrepancies between simulated and real-world behavior, known as the reality gap, challenge the transferability of test results to deployed systems. In this paper, we present a comprehensive empirical study comparing four representative testing modalities: Software-in-the-Loop (SiL), Vehicle-in-the-Loop (ViL), Mixed-Reality (MR), and full real-world testing. Using a small-scale physical vehicle equipped with real sensors (camera and LiDAR) and its digital twin, we implement each setup and evaluate two ADS architectures (modular and end-to-end) across diverse indoor driving scenarios involving real obstacles, road topologies, and indoor environments. We systematically assess the impact of each testing modality along three dimensions of the reality gap: actuation, perception, and behavioral fidelity. Our results show that while SiL and ViL setups simplify critical aspects of real-world dynamics and sensing, MR testing improves perceptual realism without compromising safety or control. Importantly, we identify the conditions under which failures do not transfer across testing modalities and isolate the underlying dimensions of the gap responsible for these discrepancies. Our findings offer actionable insights into the respective strengths and limitations of each modality and outline a path toward more robust and transferable validation of autonomous driving systems.
@@ I Introduction
Various strategies have been proposed to mitigate the reality gap. Some aim to increase simulation fidelity through high-precision modeling (digital twins [9, 10, 11]), others address specific gap dimensions, such as the perception gap, by translating simulated sensor outputs into more realistic versions using generative models [12, 13, 14, 8, 15]. However, these methods are limited to model-level testing [16, 17] and do not capture the system-level interactions between perception, planning, and control modules that govern vehicle motion. As a result, they are susceptible to actuation gaps and often miss critical system-level failures [16, 14, 18]. Other strategies involve vehicle-in-the-loop (ViL) and mixed reality (MR) testing [19], by integrating physical components such as ECUs, small-scale robots, or full vehicles into simulation loops. While ViL provides closed-loop evaluation, it remains partially virtualized and fails to capture real-world imperfections, such as sensor noise and lighting variability (perception gap) [20]. MR partially mitigates this by injecting virtual elements (e.g., obstacles) into real sensor data, enriching scenario realism. Prior system-level studies using small-scale robots investigating failure transferability [8, 21, 22, 23, 24, 25] have primarily documented the existence of the reality gap, without isolating their root causes or comparing how different test modalities influence gap reduction.
To this aim, in this paper, we conduct an empirical study of the reality gap in autonomous driving by comparing four representative testing modalities: SiL, ViL, MR, and full real-world execution (RW). Our goal is to characterize the dimensions of the reality gap, gap-specifically along actuation, perception, and behavioral fidelity, and to assess the degree to which each testing setup retains ADS behavior relative to real-world ground truth behavior. While prior work has evaluated or mitigated specific aspects of the reality gap, a broader evaluation spanning multiple testing modalities and ADS architectures remains unaddressed.
@@ III-B1 Simulator
The Donkey Car framework provides a high-fidelity Unity3D simulator that models the physical vehicle using the Nvidia PhysX engine. We extended the simulator by developing a new Unity scene that procedurally generates track layouts using Catmull-Rom splines [39], based on real-world lane margin data from Room Nominal and Generalization. To replicate the physical environments, we applied high-resolution (48 MP) floor images as textures and configured the lane markings to match each sandbox. In addition, we modified the simulator to support a virtual depth sensor that mimics the behavior of the ToF system used in our real-world setup. The virtual sensor produces depth data at the same resolution and provides intrinsic matrix parameters consistent with those retrieved by the ToF system.
@@ III-B2 Testing Framework
For the ViL and MR modalities, Figure 2 shows how the framework manages data flow between the simulator, the physical vehicle, and the ADS. Simulated data is first generated by the simulator. In the ViL setting, the full simulated sensor stream (e.g., camera or LiDAR) is used as input ❶. In the MR setting, only selected features or objects are rendered and extracted ❶(b), then merged with real sensor data from the physical vehicle ❷ to form mixed-reality inputs. The resulting sensor data, fully simulated in ViL or blended in MR-is processed by the ADS ❸, which outputs control commands for the physical vehicle ❹. The vehicle’s motion is tracked in real time ❺, and its pose is fed back to the simulator ❻ to maintain alignment between virtual and real environments. This forms a closed-loop system in which the ADS continuously perceives and acts on synchronized sensor data and vehicle dynamics.
@@ IV-D Experimental Setup Validation
For ViL and MR, the real vehicle’s motion must be continuously mapped into the simulation, requiring real-time execution to keep the simulated state and generated sensor data synchronized with the physical vehicle. Across all domains, the E2E or modular pipeline must execute fast enough to deliver control commands (simulated or real) without latency, as delays could distort the observed behavior. To verify this, we benchmarked the framework over 5000 samples, measuring execution time for every component. All modalities (SiL, ViL, MR, RW) ran in real time: the simulator achieved $\approx 80$ FPS ($\approx 12.5$ ms per frame) on our workstation (AMD Ryzen 5, 16 GB RAM, NVIDIA RTX 2060). The tracking system has been measured (at the simulator-level) to provide state updates at $\approx 99.93$ Hz ($\approx 10.01$ ms). Since the onboard camera and LiDAR sensors operate at 20 FPS (50 ms per frame), we configure the simulator to run its camera at 20 FPS, and to avoid unnecessary system load, we cap the global simulation frame-rate at 60 FPS. The E2E ADS executed in $\approx 15$ ms on average, and the modular ADAS in $\approx 20$ ms. Inference begins as soon as a new sensor message arrives and can run in parallel if needed, though both pipelines consistently complete before the next frame. Control commands, measured at the vehicle-level, were issued every $\approx 50.5$ ms, remaining stable across all experiments.
@@ IV-E2 RQ2 (actuation gap)
To address RQ2, we quantify the actuation gap by comparing the physical response of the real vehicle to that of a simulated vehicle under identical control inputs. We limit the evaluation to SiL, ViL, and RW, omitting MR as this modality does not change actuation fidelity beyond what is already captured by ViL. Each experiment was repeated five times to capture variability, and data were logged at 100 Hz, yielding large frame-level sample sizes. Statistical analysis is applied only to per-frame metrics, where sample sizes support meaningful significance testing. We design five test scenarios spanning both low-level actuation and high-level control, described next. Forward Motion. Constant throttle is applied with steering fixed at zero and braking disabled. Throttle commands are taken directly from the RW-recorded ROS messages with matched timestamps. Trials run for three seconds using throttle values of 0.34, 0.365, and 0.39, which represent the lowest throttle at which the vehicle starts moving, to the highest throttle at which the vehicle can experiment while remaining in the tracking area. We compute the total distance traveled, average speed, and trajectory, and compare both speed and trajectory distributions. This setup results in 15 executions (3 speed values × 5 repetitions).
@@ V-D1 Internal validity
We compared all ADS under identical parameter settings. One threat to internal validity concerns our custom implementation. However, this was unavoidable as no similar evaluation frameworks are available, to the best of our knowledge. Another threat may be due to our data collection phase and training of ADS, which may exhibit a large number of misbehaviors if trained inadequately or with poor-quality data. We mitigated this threat by training and fine-tuning the best publicly available driving models, which performed consistently in nominal RW conditions. The RW-trained E2E model completed scenarios N1 and N2 in all five RW runs without failure (Table 1, Nominal–E2E–Real), indicating adequate training. The LiDAR-based modular pipeline failed in N2 due to late-appearing obstacles, as its perception relies on deterministic clustering rather than learned models; however, the ablation study shows the planning module operates correctly when decoupled from LiDAR inputs. Failures in scenario G stem from domain shift rather than poor training: the E2E ADS succeeds in SiL but fails in RW, which would not occur if the model were fundamentally incapable of solving the task. Overall, these results confirm that our comparisons are not confounded by under-trained or low-quality models.
@@ V-D2 External validity
We used a limited number of ADS architecture models in our evaluation, which we mitigated by covering representative ADS architectures. We considered only two physical tracks and a scaled-down platform, which may not capture all real-world physics (e.g., suspension dynamics or vehicle mass distribution). However, our goal is not to model absolute performance but to assess relative fidelity across modalities. Donkey Car was used as a proxy for full-size ADS also in previous studies [55, 21, 22, 23, 8, 14] and uniquely satisfies our requirements for studying transferability between simulated and real-world testing of ADS. Other platforms, such as DeepPiCar [56] and JetRacer [57], lack integrated simulators; Roboracer [58] offers only low-fidelity physical simulations that do not realistically capture real-world driving dynamics. AWS DeepRacer [59] is tightly integrated with AWS infrastructure and is primarily designed for reinforcement learning use cases, which are outside the scope of this work. In contrast, Donkey Car has been successfully adopted in numerous real-vehicle autonomous driving studies [55, 21, 22, 23, 8, 14, 36, 37, 38], making it a practical and cost-effective experimentation platform.
@@ VI-A Dissecting and Addressing the ADS Reality Gap
To mitigate the former in SiL, without relying on physical vehicles, recent neural rendering approaches have been proposed. Examples include generative AI for translating simulated into photorealistic images [12, 8, 14, 60, 13, 15] and diffusion models for generating realistic operational design domains [61, 62, 63]. Some are already integrated in simulators such as NVIDIA Omniverse. While promising, these methods still suffer from correctness issues (e.g., artifacts, hallucinations), increase runtime cost, and leave the actuation gap and simulator unreliability [64, 65] unaddressed.
@@ VII-A Reality Gap Assessment Studies
The reality gap has been subjected to active research in many fields, including robotics, automotive, and artificial intelligence. For a comprehensive survey, we refer the reader to Hu et al. [75], and contributions in the software engineering community [76, 1, 64, 8, 2, 3]. Concerning empirical studies on the reality gap, Stocco et al. [8] compare ADS lane-keeping models in simulated and physical environments, highlighting critical shortcomings that contribute to the gap. In this work, beyond assessing the gap between SiL and RW, we also evaluate mitigation strategies such as ViL and MR. Similarly, Gao et al.[77] propose MultiTest, a physically-aware object insertion framework for testing the robustness of fusion-based perception systems, while Gao et al. [78] outline key challenges in benchmarking AI-enabled multi-sensor fusion across diverse conditions. However, prior studies focus mainly on perception robustness in simulation and stop short of full-system evaluation. To our knowledge, our work offers the first unified, system-level assessment of ViL and MR as reality-gap mitigation strategies, analyzing their impact on perception, actuation, and behavior during live execution in SiL, MR, and RW under consistent, controlled conditions. However, the transferability to real vehicles is not assessed.
@@ VII-B Reality Gap Mitigation Studies
To address the perception gap, researchers have explored the use of generative AI to translate simulated data into photorealistic images [12, 8, 14, 13, 15], LiDAR point clouds [80], and methods for synthesizing operational design domains with high visual fidelity [61, 62, 81, 82]. While existing methods enhance perceptual realism, they typically run offline at the model level or in SiL, targeting single sensor modalities. In contrast, our work evaluates perception fidelity during live, system-level execution across SiL, MR, and RW, directly comparing how perceptual gaps affect the ADS behavior.